The ratio between the root mean square speed of H2 at 50K and that of O2 at 800 K is
a. 4
b. 2
c. 1
d. 1/4

Grade:upto college level

2 Answers

Jitender Pal
9 years ago
Sol. Expression of rms is : urms = √(3RT/M) ⇒ (u_rms (H_2 at 50 K))/(u_rms (O_2 at 800 K) ) = √((3R ×50)/2)/√((3R ×800)/32) = √(50/2 × 32/800) = 1
Rishi Sharma
3 years ago
Dear Student,
Please find below the solution to your problem.

Expression of rms is : urms = √(3RT/M)
⇒ (u_rms (H_2 at 50 K))/(u_rms (O_2 at 800 K) ) = √((3R ×50)/2)/√((3R ×800)/32)
= √(50/2 × 32/800)
= 1

Thanks and Regards